A balanced, positive-sequence, wye-connected source has V- 220/0 V rms and supplies an unbalanced wye-connected load ZAN-7 10 Ω, ZBN-8 +j6 Ω, and ZCN 10-j12 Ω via transmission lines with impedances ZA-2 +jl Ω, ZB-2-j0.5 Ω, and ZC-2 +j | Ω (a) Label all node voltages and loop currents (b) Determine the currents IaA? IbB, and 1cC. (c) Determine the voltages VAN, N, and VCN (d) Determine the voltages VAR, VRc, and VCA (e) (i) Determine the total complex power absorbed by the load (ii) Determine the total complex power of the source (f) Determine the total complex power losses in the transmission line 220/0° V( 220-120° V AN ZB 220/120 V BN C. Z Fig. P1.1

Explain Please.
Answer #1

ZAN 2CN Apply kvL fo (cop 0-330-JI9o5 -j39/ ouO3.2s - jiu3u -19-j16 S lo-s != 20.27 -19-j/6.5-330- -J lo.le =e2.6s I.Lk53 2- -38 ー853 8.75 +j 3S 19 -19-= 262,6ャ65) -310? -j366 53 vefr = VCN-YnN 。 (7tJ3//.24)一643.0,+513,.86) Total comples pucet erpe abserbedl by the loe is CN C